A tire factory wants to set a minimum mileage guarantee on its tyre. Tyre tests reveals that the mean mileage has a normal distribution with mean 109 and standard deviation 2.65.
the manufacturer wants to set the minimum guaranteed mileage so that no more than 2.5 percent of tyres will have to be replaced.
The Lower limit of guaranteed mileage is :?

Answer by Boreal(15235)

z(0.025)=-1.96
z=(x-mean)/sd
-1.96=(x-109)/2.65
-5.194=x-109
x=103.81